If two dice are rolled one time, find the probability of getting these results:
a. A sum of 5
b. A sum of 9 or 10
c. Doubles
1
Expert's answer
2020-02-03T09:22:21-0500

a) P(2,3 or 3,2 or 4,1 or 1,4)

=P(2,3)+ P(3,2)+P(4,1)+P(4,1)

"=\frac{1}{36}+\frac{1}{36}+\frac{1}{36}+\frac{1}{36}=\frac{4}{36}=\frac{1}{9}"

b) P(5,4 or 4,5 or 6,3 or 3,6 or 5,5 or 4,6 or 6,4)

"=\frac{1}{36}\times{7}=\frac{7}{36}"

c) P(1,1 or 2,2 or 3,3 or 4,4 or 5,5 or 6,6)

"=\frac{1}{36}\times{6}=\frac{1}{6}"